You noticed that the exchange rate between the Korean won and the U.S. dollar has changed considerably. The won/dollar...

  1. You noticed that the exchange rate between the Korean won and the U.S. dollar has changed considerably. The won/dollar exchange rate has moved from 1200 won per dollar to 800 won per dollar. By what percentage has the value of the dollar changed with respect to the won? Show your work. (4 points)
Answer #1

% change in the dollar value = [Value now - Value earlier] / Value earlier

= [800 - 1000] / 1000 = -200 / 1000 = -0.20, or -20%